OpenStudy (anonymous):

Given f(x) = the quantity of x plus 4, divided by 6, solve for f-1(7).

OpenStudy (anonymous):

\[f(x)=\frac{ x+4 }{ 6 }\]

OpenStudy (anonymous):

solve for f^-1(7).

OpenStudy (anonymous):

@Euler271 @amistre64 @agent0smith

OpenStudy (anonymous):

hello? o.o

OpenStudy (anonymous):

@cwrw238

OpenStudy (agent0smith):

Can you find the inverse function? solve this for x (rearrange to get x=...) \[\large y=\frac{ x+4 }{ 6 }\]

OpenStudy (anonymous):

x(y)=2(3y-2) ?

OpenStudy (anonymous):

@helpme1.2 this one.

OpenStudy (agent0smith):

You didn't need to factor it at all... x = 6y - 4 is fine. Now swap x and y. SO f^-1(x) = 6x - 4

OpenStudy (agent0smith):

And find r f^-1(7)

OpenStudy (anonymous):

38...?

OpenStudy (anonymous):

i'm guessing that's it? xD

OpenStudy (agent0smith):

if 6*7 - 4 is 38 then yes that's it :P